# Spatial4D-Bench: A Versatile 4D Spatial Intelligence Benchmark

### v1

@@ 3.3 Task Taxonomy
To systematically evaluate the gap between current MLLMs and human-level spatial cognition, as shown in Figure˜4, we organize the 18 tasks into a hierarchical taxonomy composed of six core categories: object understanding, scene understanding, spatial relationship understanding, spatiotemporal relationship understanding, spatial reasoning and spatiotemporal reasoning. This taxonomy progresses from object/scene-level perception, through spatial/spatiotemporal understanding, to dynamic spatial/spatiotemporal reasoning, mirroring the cognitive abilities of human intelligence [waller2013handbook, newcombe2004spatial].
@@
Spatial intelligence begins with the accurate perception of intrinsic physical properties, which is a prerequisite for interacting with a physical scene. This category includes object size estimation, object attribute estimation, object counting and affordance estimation. A robust spatial agent must ground visual features into both precise metric values (size) and semantic properties (attributes). While existing benchmarks such as VSI-Bench [Yang_2025_CVPR] evaluate metric properties, they often overlook intrinsic attributes such as color, shape, and material. Spatial4D-Bench integrates both metric and attribute understanding into a unified evaluation framework. Crucially, we also incorporate affordance estimation here as a functional attribute. Unlike SIBench [yu2025far] or SpaCE-10 [gong2025space], which focus on simple existence queries (e.g., “is there an item to sit on?”), we integrate affordance with 3D grounding, requiring the model to identify the specific spatial position that satisfies a functional description (e.g., “which position allows me to wash hands?”). This verifies that MLLMs possess a comprehensive understanding of an object’s physical characteristics, which is a fundamental prerequisite for complex interaction tasks.
@@
Beyond isolated entities, an agent must comprehend the global semantic and geometric context of the environment to answer “what” is present and “where” it belongs. This category encompasses room size estimation, scene classification, and 3D grounding. Here, the model must abstract local visual cues into a broader understanding of the environment type and its volumetric scale. Crucially, we include 3D grounding (object detection and localization in 3D space) in this category as it represents the population of the scene layout. Unlike benchmarks like InternSpatial [deng2025internspatial] that rely on 2D bounding boxes, our 3D grounding task requires 3D spatial grounding, ensuring that MLLMs possess a volumetric understanding of object placement.
@@
Once entities and the scene are defined, a spatial intelligence model needs to be able to resolve the geometric topology and metric layout between them. This category is composed of relative distance estimation, absolute distance estimation and relative orientation estimation. Navigating the physical world requires resolving spatial ambiguities. By evaluating both relative (topological) and absolute (metric) distances, we test the robustness of the model’s depth perception. While benchmarks like SpatialRGPT [cheng2024spatialrgpt] address relative spatial relations, we enforce a more strict evaluation by combining relative topology with absolute metric estimation, ensuring that the model is not merely guessing based on 2D perspective cues.
@@
Incorporating the fourth dimension of time, this category evaluates the ability to track object states and dynamics over time. This category represents the leap from 3D to 4D spatial intelligence. Tasks include action recognition, appearance order, spatial memory, and state change detection. The real world is a dynamic environment in which spatial and temporal aspects continuously evolve, and humans living in such an environment inherently develop 4D spatial intelligence that supports a wide range of spatial understanding and reasoning abilities. While VSI-Bench includes basic appearance ordering [Yang_2025_CVPR] where clearly visible objects need to be identified in terms of the sequence in which they show up in a video, true 4D intelligence requires maintaining a coherent world model of objects even when they are occluded. Our spatial memory task evaluates this working memory capability by requiring the models to track objects that exit the field of view. Action recognition evaluates the model’s ability to semantically categorize dynamic events within 3D space, bridging the gap between object detection and dynamic scene understanding. Furthermore, state change detection tests the causal understanding of interactions over time, such as a door opening rather than just movement tracking, distinguishing between simple movement and meaningful state transitions (e.g., a door opening).
@@
Moving from passive perception to active embodiment, this category assesses the agent’s ability to reason about its own state and plan movements. It is composed of egocentric reasoning and route plan. We introduce egocentric reasoning to challenge the model to infer the observer’s own spatial state (e.g., “how did the camera rotate?”), which is a prerequisite for self-localization. This capability is critical for embodied agents operating in first-person views to localize themselves within the environment, moving beyond purely allocentric reasoning. Building on this, our route plan task elevates the difficulty beyond the single-room queries of VSI-Bench [Yang_2025_CVPR]. Instead of using single-room single-step queries found in previous benchmarks [Yang_2025_CVPR], we require long-horizon trajectory planning across multiple rooms, simulating the sequential cognitive load of a mobile robot operating in a complex environment. To answer correctly, the model must predict a valid sequence of actions (e.g., selecting the correct directions for multiple blanks in an instruction set) rather than a single directional move.
@@ 4.1 Evaluation Setup
We benchmark a diverse set of state-of-the-art (SOTA) multimodal large language models, comprising both leading proprietary models (GPT-5 [OpenAI2025] and Gemini 2.5-Pro [Gemini2025]) and top-tier open-source models (VideoLLama3 [zhang2025videollama], Qwen2.5-VL [bai2025qwen2], Qwen3-VL [bai2025qwen2], and InternVL3.5 [wang2025internvl3]). To analyze the impact of model capacity, we evaluate variants across model sizes ranging from 7B to 241B parameters. In total, we comprehensively evaluate eleven models on Spatial4D-Bench. At the time of evaluation, these models represented the state-of-the-art. We note that the results reported for proprietary models reflect the versions accessible during our experimental window; as these API-based services are subject to continuous updates by their providers, exact reproducibility may vary over time. For memory-intensive models (e.g., 241B parameters), we adopt a uniform sampling strategy of $64$ frames from each video to mitigate memory constraints during evaluation following [Yang_2025_CVPR].
@@
We conduct all evaluations under a zero-shot setting. Adhering to the protocol established in [Yang_2025_CVPR], we compute the metrics tailored to the distinct answer formats in our Spatial4D-Bench, including Multiple-Choice Answers (MCA) and Numerical Answers (NA). For MCA tasks, we report exact matching with possible fuzzy matching. For NA tasks, we evaluate performance using the Mean Relative Accuracy (MRA). The MRA measures the model’s consistency across a spectrum of error tolerances, calculated as the average satisfaction rate over a set of thresholds $\mathcal{C}$, which is defined as follows:
@@
Similar to [Yang_2025_CVPR], to contextualize MLLM performance, we compare against the baselines of: Human Level Performance and Chance Level Baselines (Random and Frequency). For Human Level Performance, we sample 1,000 QA pairs from Spatial4D-Bench as a representative subset and recruited several qualified human evaluators with relevant research background to independently complete the test. To establish a theoretical ceiling for the performance of human-level 4D spatial intelligence, we report the highest score among human evaluators. For Chance Level (Random), we report the expected average accuracy of random selection (MCA only). For Chance Level (Frequency), we report performance by always selecting the dataset’s most frequent answer for each task. This indicates the potential performance gains that could be obtained due to the long-tail distribution of answers or the imbalance in multiple-choice distributions.
@@ 4.2 Main Evaluation Results
We present the main evaluation results in Table˜2 and show radar chart visualization of model performance across 18 tasks in Figure˜5. Our analysis and discussion of the main evaluation results are as follows.
@@
From Table˜2 and Figure˜5(a), we can see that the overall 4D spatial intelligence performance of MLLMs remains significantly inferior to that of humans. The best proprietary model, GPT-5, achieves an average score of $60.90$, while the best open-source model, Qwen3-VL-235B-A22B, yields an average score of $56.17$. This indicates that the performance gap between the best proprietary MLLMs and the best open-source MLLMs is relatively small, although the best proprietary MLLMs still perform better. However, the overall performance gap between MLLMs (both proprietary and open-source) and human-level 4D spatial intelligence remains significant. On Spatial4D-Bench, humans achieve an average score of $78.02$, outperforming the best proprietary model (GPT-5) by approximately $17$ and the best open-source model (Qwen3-VL-235B-A22B) by $22$. Current MLLMs appear to operate as “frame-based observers” rather than “world observers”. They lack an intuitive physics engine or coherent temporal memory, preventing effective reasoning about causality, permanence, and dynamics. Spatial4D-Bench effectively exposes this 4D reasoning gap, which was invisible in existing benchmarks. Substantial efforts are required to further improve MLLMs toward human-level spatial cognition.
@@
As shown in Table˜2, Figure˜5(b) and Figure˜5(c), for the fundamental object/scene understanding category (e.g.  object size, object counting, and room size), existing MLLMs often reach or even surpass human performance. In particular, Qwen3-VL-30B-A3B achieves $80.10$ on the object size estimation task, outperforming the human baseline of $74.61$. In addition, on the object counting task, the best MLLM also outperforms human performance. We attribute this to distinct cognitive processing differences. First, while humans rely on intuitive relative scale, they struggle with precise absolute metric estimation (e.g., “is this table 1.2m or 1.4m?”) from 2D projections without explicit reference scales while MLLMs leverage massive prior knowledge obtained by pre-training on 3D and geometric data which are useful. Moreover, it is noticed that external factors can also result in low accuracy of the human evaluation. For example, in object counting, to elevate the the difficulty level in model evaluation, Spatial4D-Bench includes a set of low-resolution videos with significant jittering and inconsistent frames compared with videos used in [Yang_2025_CVPR], posing a huge challenge to human evaluators in correctly matching and identifying the objects within these videos. We have also found that discrepancies among humans in defining the object categories also play an important role in the low accuracy of this task. While a nightstand is considered to be a table in the QA design, a human evaluator may not count it as a table, causing inconsistent result with the ground truth answer.
@@
The primary source of the human-AI gap lies in reasoning-related tasks. As shown in Table˜2, Figure˜5(e), and Figure˜5(f), MLLMs exhibit substantial performance degradation on the spatial reasoning and spatiotemporal reasoning categories. These challenging categories are the core of Spatial4D-Bench. Specifically, for spatiotemporal reasoning, even the top-tier MLLMs lag significantly behind humans (e.g., GPT-5 scores $32.83$ on route plan vs. Human $91.67$). This $60\%$ gap highlights a fundamental deficiency of MLLMs in maintaining a coherent 4D world model over extended temporal sequences. For the physical plausibility reasoning task, MLLMs score near random chance ($30\%-40\%$), whereas humans intuitively reject these physically impossible scenarios ($66.67\%$). This indicates that SOTA MLLMs struggle to ground visual perception in fundamental physical laws.
@@
Table 3: Comparison with VSI-Bench [Yang_2025_CVPR] on overlapping tasks. Overall, state-of-the-art models perform worse on our Spatial4D-Bench, especially on the redefined route plan task. These results indicate that our Spatial4D-Bench is more challenging than VSI-Bench [Yang_2025_CVPR] on overlapping tasks.
@@ 4.3 Benchmark Challenge Analysis
As summarized in Table˜1, Spatial4D-Bench provides a more comprehensive evaluation benchmarks with 18 tasks. Since there are some tasks overlapping with existing benchmarks, it would be interesting to investigate the challenge of these tasks between existing benchmarks and Spatial4D-Bench. To this end, we compare with VSI-Bench [Yang_2025_CVPR], one of the most representative spatial intelligence benchmarks, on overlapping tasks using Qwen3-VL-30B-A3B, InternVL3.5-38B, and InternVL3.5-8B. As shown in Table˜3, overall, state-of-the-art models perform worse on our Spatial4D-Bench. Specifically, on some perception-related tasks, e.g., object size and room size, the performance is comparable across both benchmarks. However, for more challenging reasoning tasks, such as relative orientation and route plan, substantial divergences can be observed, especially on the route plan task, where the tested models exhibit a substantial performance drop of approximately 22% to 29% on Spatial4D-Bench compared to VSI-Bench [Yang_2025_CVPR]. In addition to the difficulty elevation described in Section˜3.3, this can be attributed to the fact that the route plan task in Spatial4D-Bench requires long-horizon planning and 4D spatial understanding over longer sequences and more complex state transitions, which distinguishes Spatial4D-Bench from existing benchmarks that may rely on shorter horizons or simpler topological graphs.
@@ 4.4 Further Analysis and Discussion
      Language Priors vs. Random Chance: In the absence of visual data, the model defaults to learned LLM priors, e.g., probabilistically associating “oven” with “kitchen”. As similarly observed on VSI-Bench [Yang_2025_CVPR], these priors often provide a statistical floor that is significantly better than random chance.
@@ 4.5 Qualitative Case Analysis
A major challenge in 4D spatial intelligence is spatiotemporal reasoning that requires understanding a coherent 4D world over time rather than treating videos as a collection of disjoint semantic concepts. By designing tasks that connect past, present, and future visual observations, Spatial4D-Bench reveals a significant temporal incoherence in SOTA MLLMs. Figure˜6 shows failure and success cases from the action prediction task, which exposes this fragility. In Case 1 (failure), the ground truth action is “washing” a yogurt cup, but the model incorrectly predicts “placing it in the sink”, driven by a semantic prior rather than the specific temporal action. The benchmark’s design allows us to probe the cause via the model’s self-explanation: “…the person walks to the sink holding the empty cup… indicating they are about to place it in the sink rather than wash it or throw it away”. This textual feedback from the model demonstrates that the model successfully tracked the trajectory but failed to infer the latent intent (washing for recycling), proving that even powerful models fail to ground their predictions in the actual temporal dynamics when those dynamics conflict with training priors. On the other hand, Case 2 (success) validates the benchmark’s ability to measure robust chain-of-thought reasoning when explicit cues (OCR) are present. The model correctly infers the drink will taste “sweet” by synthesizing the label “sugar” with the stirring action (“mixing sugar into the drink makes it sweet…”). This contrast underscores the diagnostic value of our Spatial4D-Bench: it can distinguish between scenarios where models rely on robust textual grounding versus those where they collapse into hallucination due to temporal ambiguity.
@@
By incorporating the physical plausibility reasoning in the spatiotemporal reasoning category, specifically using AI-generated physical anomalies, Spatial4D-Bench provides a unique evaluation of the “intuitive physics engine” of MLLMs. Our evaluation exposes a sharp dissociation between low-level perceptual grounding and high-level physical knowledge. In Figure˜7 Case 1 (failure), the benchmark challenges the model with a subtle violation of fluid dynamics. GPT-5 fails to identify a physical anomaly where liquid levels behave inconsistently during pouring, yet its self-explanation recites perfect high-level physics principles: “Matter doesn’t disappear from one place and reappear somewhere else… Teleportation breaks locality and conservation of mass…”. This result highlights a critical insight enabled by our benchmark: SOTA models possess abstract knowledge of physical laws but lack the visual grounding to detect their violation in pixel space. The model “knows” the laws of physics but cannot “see” them being broken, relying instead on analyzing the textual plausibility of the options (e.g., rejecting “teleportation” as a concept) rather than verifying the visual dynamics. Spatial4D-Bench thus serves as a necessary filter to differentiate between models that merely know physics textually and those that can perceive physics visually. In contrast, Case 2 (success) shows that the model is capable of detecting violations in rigid body dynamics. It correctly identifies that a baseball and bat exhibit unnatural softness and defy gravity, suggesting that conspicuous deviations in material properties and trajectory dynamics are easier for current architectures to flag than subtle fluid inconsistencies.
@@
Through evaluation on Spatial4D-Bench, we observe that perceptual ambiguity can lead to confident hallucinations, even in tasks (such as spatial memory) that require explicit tracking. In Figure˜8 Case 1 (failure), the model incorrectly localizes the final position of a teapot. Crucially, the self-explanation reveals a confident hallucination: “After he pours the water… he places the teapot on the back-right burner (the burner directly above the original position)”. This contradicts the visual evidence of the teapot’s placement in other locations. This suggests that texture similarity across the stove top surface causes the model to lose track of the object’s specific geometric coordinates. Unlike a tracking failure where a model might express uncertainty, here the model constructs a coherent (but false) narrative to fill the perceptual gap. This failure mode validates the necessity of Spatial4D-Bench’s fine-grained annotation: unlike simpler existence or classification tasks, our spatial memory queries force the model to confront texture confusion and occlusion. The benchmark demonstrates that even when a model is confident and generates plausible-sounding narratives, it often lacks the precise metric grounding required for 4D spatial intelligence. By contrast, Case 2 (success) demonstrates that the model maintains robust temporal tracking when visual events are semantically distinct. The model correctly counts that the faucet was used “two” times. Its self-explanation, “…first to wash their hands at the sink, and later to fill a glass with water”, shows a successful linking of two separate temporal events, implying that the “grounding gap” is highly sensitive to visual saliency. The model fails when tracking requires resolving low-level texture ambiguity (Case 1) but succeeds when tracking high-level, semantically distinct actions (Case 2), further validating the benchmark’s ability to probe the granular limits of spatial memory.
@@
The route plan task in the spatial reasoning category illustrates the inability of current MLLMs to construct accurate mental maps from egocentric video streams. In Figure˜9, the model attempts to plan a path from a hallway to a bathroom. While GPT-5 outputs a confident self-explanation, reasoning that a right turn at step 7 would “line up with the bedroom doorway”, this contradicts the visual geometry: the robot has just entered a hallway configuration where a left turn (option A) is geometrically required to face the bedroom entrance. The model hallucinates a spatial layout that fits its internal narrative but ignores the visual reality of the scene. Furthermore, while the model correctly identifies the final turn (step 11) as a right turn, this success is coincidental, derived from a flawed intermediate trajectory. This failure highlights a critical bottleneck: SOTA MLLMs struggle with long-horizon spatial consistency and relative orientation. By enforcing strict directional accuracy over multi-step paths, Spatial4D-Bench effectively disentangles true embodied route planning capabilities from lucky guesses driven by language priors.

### v2

@@ 3.3 Task Taxonomy
To systematically evaluate the gap between current MLLMs and human-level spatial cognition, as shown in Fig.˜4, we organize the 18 tasks into a hierarchical taxonomy composed of six core categories: object understanding, scene understanding, spatial relationship understanding, spatiotemporal relationship understanding, spatial reasoning and spatiotemporal reasoning. This taxonomy progresses from object/scene-level perception, through spatial/spatiotemporal understanding, to dynamic spatial/spatiotemporal reasoning, mirroring the cognitive abilities of human intelligence [35, 25].
@@
Spatial intelligence begins with the accurate perception of intrinsic physical properties, which is a prerequisite for interacting with a physical scene. This category includes object size estimation, object attribute estimation, object counting and affordance estimation. A robust spatial agent must ground visual features into both precise metric values (size) and semantic properties (attributes). While existing benchmarks such as VSI-Bench [42] evaluate metric properties, they often overlook intrinsic attributes such as color, shape, and material. Spatial4D-Bench integrates both metric and attribute understanding into a unified evaluation framework. Crucially, we also incorporate affordance estimation here as a functional attribute. Unlike SIBench [46] or SpaCE-10 [17], which focus on simple existence queries (e.g., “is there an item to sit on?”), we integrate affordance with 3D grounding, requiring the model to identify the specific spatial position that satisfies a functional description (e.g., “which position allows me to wash hands?”). This verifies that MLLMs possess a comprehensive understanding of an object’s physical characteristics, which is a fundamental prerequisite for complex interaction tasks.
@@
Beyond isolated entities, an agent must comprehend the global semantic and geometric context of the environment to answer “what” is present and “where” it belongs. This category encompasses room size estimation, scene classification, and 3D grounding. Here, the model must abstract local visual cues into a broader understanding of the environment type and its volumetric scale. Crucially, we include 3D grounding (object detection and localization in 3D space) in this category as it represents the population of the scene layout. Unlike benchmarks like InternSpatial [13] that rely on 2D bounding boxes, our 3D grounding task requires 3D spatial grounding, ensuring that MLLMs possess a volumetric understanding of object placement.
@@
Once entities and the scene are defined, a spatial intelligence model needs to be able to resolve the geometric topology and metric layout between them. This category is composed of relative distance estimation, absolute distance estimation and relative orientation estimation. Navigating the physical world requires resolving spatial ambiguities. By evaluating both relative (topological) and absolute (metric) distances, we test the robustness of the model’s depth perception. While benchmarks like SpatialRGPT [9] address relative spatial relations, we enforce a more strict evaluation by combining relative topology with absolute metric estimation, ensuring that the model is not merely guessing based on 2D perspective cues.
@@
Incorporating the fourth dimension of time, this category evaluates the ability to track object states and dynamics over time. This category represents the leap from 3D to 4D spatial intelligence. Tasks include action recognition, appearance order, spatial memory, and state change detection. The real world is a dynamic environment in which spatial and temporal aspects continuously evolve, and humans living in such an environment inherently develop 4D spatial intelligence that supports a wide range of spatial understanding and reasoning abilities. While VSI-Bench includes basic appearance ordering [42] where clearly visible objects need to be identified in terms of the sequence in which they show up in a video, true 4D intelligence requires maintaining a coherent world model of objects even when they are occluded. Our spatial memory task evaluates this working memory capability by requiring the models to track objects that exit the field of view. Action recognition evaluates the model’s ability to semantically categorize dynamic events within 3D space, bridging the gap between object detection and dynamic scene understanding. Furthermore, state change detection tests the causal understanding of interactions over time, such as a door opening rather than just movement tracking, distinguishing between simple movement and meaningful state transitions (e.g., a door opening).
@@
Moving from passive perception to active embodiment, this category assesses the agent’s ability to reason about its own state and plan movements. It is composed of egocentric reasoning and route plan. We introduce egocentric reasoning to challenge the model to infer the observer’s own spatial state (e.g., “how did the camera rotate?”), which is a prerequisite for self-localization. This capability is critical for embodied agents operating in first-person views to localize themselves within the environment, moving beyond purely allocentric reasoning. Building on this, our route plan task elevates the difficulty beyond the single-room queries of VSI-Bench [42]. Instead of using single-room single-step queries found in previous benchmarks [42], we require long-horizon trajectory planning across multiple rooms, simulating the sequential cognitive load of a mobile robot operating in a complex environment. To answer correctly, the model must predict a valid sequence of actions (e.g., selecting the correct directions for multiple blanks in an instruction set) rather than a single directional move.
@@ 4.1 Evaluation Setup
We benchmark a diverse set of state-of-the-art (SOTA) multimodal large language models, comprising both leading proprietary models (GPT-5 [26] and Gemini 2.5-Pro [10]) and top-tier open-source models (VideoLLama3 [47], Qwen2.5-VL [41], Qwen3-VL [41], and InternVL3.5 [36]). To analyze the impact of model capacity, we evaluate variants across model sizes ranging from 7B to 241B parameters. In total, we comprehensively evaluate eleven models on Spatial4D-Bench. At the time of evaluation, these models represented the state-of-the-art. We note that the results reported for proprietary models reflect the versions accessible during our experimental window; as these API-based services are subject to continuous updates by their providers, exact reproducibility may vary over time. For memory-intensive models (e.g., 241B parameters), we adopt a uniform sampling strategy of $64$ frames from each video to mitigate memory constraints during evaluation following [42].
@@
We conduct all evaluations under a zero-shot setting. Adhering to the protocol established in [42], we compute the metrics tailored to the distinct answer formats in our Spatial4D-Bench, including Multiple-Choice Answers (MCA) and Numerical Answers (NA). For MCA tasks, we report exact matching with possible fuzzy matching. For NA tasks, we evaluate performance using the Mean Relative Accuracy (MRA). The MRA measures the model’s consistency across a spectrum of error tolerances, calculated as the average satisfaction rate over a set of thresholds $\mathcal{C}$, which is defined as follows:
@@
Similar to [42], to contextualize MLLM performance, we compare against the baselines of: Human Level Performance and Chance Level Baselines (Random and Frequency). For Human Level Performance, we sample 1,000 QA pairs from Spatial4D-Bench as a representative subset and recruited several qualified human evaluators with relevant research background to independently complete the test. To establish a theoretical ceiling for the performance of human-level 4D spatial intelligence, we report the highest score among human evaluators. For Chance Level (Random), we report the expected average accuracy of random selection (MCA only). For Chance Level (Frequency), we report performance by always selecting the dataset’s most frequent answer for each task. This indicates the potential performance gains that could be obtained due to the long-tail distribution of answers or the imbalance in multiple-choice distributions.
@@ 4.2 Main Evaluation Results
We present the main evaluation results in Table˜2 and show radar chart visualization of model performance across 18 tasks in Fig.˜5. Our analysis and discussion of the main evaluation results are as follows.
@@
From Table˜2 and Fig.˜5(a), we can see that the overall 4D spatial intelligence performance of MLLMs remains significantly inferior to that of humans. The best proprietary model, GPT-5, achieves an average score of $60.90$, while the best open-source model, Qwen3-VL-235B-A22B, yields an average score of $56.17$. This indicates that the performance gap between the best proprietary MLLMs and the best open-source MLLMs is relatively small, although the best proprietary MLLMs still perform better. However, the overall performance gap between MLLMs (both proprietary and open-source) and human-level 4D spatial intelligence remains significant. On Spatial4D-Bench, humans achieve an average score of $78.02$, outperforming the best proprietary model (GPT-5) by approximately $17$ and the best open-source model (Qwen3-VL-235B-A22B) by $22$. Current MLLMs appear to operate as “frame-based observers” rather than “world observers”. They lack an intuitive physics engine or coherent temporal memory, preventing effective reasoning about causality, permanence, and dynamics. Spatial4D-Bench effectively exposes this 4D reasoning gap, which was invisible in existing benchmarks. Substantial efforts are required to further improve MLLMs toward human-level spatial cognition.
@@
As shown in Table˜2, Fig.˜5(b) and Fig.˜5(c), for the fundamental object/scene understanding category (e.g.  object size, object counting, and room size), existing MLLMs often reach or even surpass human performance. In particular, Qwen3-VL-30B-A3B achieves $80.10$ on the object size estimation task, outperforming the human baseline of $74.61$. In addition, on the object counting task, the best MLLM also outperforms human performance. We attribute this to distinct cognitive processing differences. First, while humans rely on intuitive relative scale, they struggle with precise absolute metric estimation (e.g., “is this table 1.2m or 1.4m?”) from 2D projections without explicit reference scales while MLLMs leverage massive prior knowledge obtained by pre-training on 3D and geometric data which are useful. Moreover, it is noticed that external factors can also result in low accuracy of the human evaluation. For example, in object counting, to elevate the the difficulty level in model evaluation, Spatial4D-Bench includes a set of low-resolution videos with significant jittering and inconsistent frames compared with videos used in [42], posing a huge challenge to human evaluators in correctly matching and identifying the objects within these videos. We have also found that discrepancies among humans in defining the object categories also play an important role in the low accuracy of this task. While a nightstand is considered to be a table in the QA design, a human evaluator may not count it as a table, causing inconsistent result with the ground truth answer.
@@
The primary source of the human-AI gap lies in reasoning-related tasks. As shown in Table˜2, Fig.˜5(e), and Fig.˜5(f), MLLMs exhibit substantial performance degradation on the spatial reasoning and spatiotemporal reasoning categories. These challenging categories are the core of Spatial4D-Bench. Specifically, for spatiotemporal reasoning, even the top-tier MLLMs lag significantly behind humans (e.g., GPT-5 scores $32.83$ on route plan vs. Human $91.67$). This $60\%$ gap highlights a fundamental deficiency of MLLMs in maintaining a coherent 4D world model over extended temporal sequences. For the physical plausibility reasoning task, MLLMs score near random chance ($30\%-40\%$), whereas humans intuitively reject these physically impossible scenarios ($66.67\%$). This indicates that SOTA MLLMs struggle to ground visual perception in fundamental physical laws.
@@
Table 3: Comparison with VSI-Bench [42] on overlapping tasks. Overall, state-of-the-art models perform worse on our Spatial4D-Bench, especially on the redefined route plan task. These results indicate that our Spatial4D-Bench is more challenging than VSI-Bench [42] on overlapping tasks.
@@ 4.3 Benchmark Challenge Analysis
As summarized in Table˜1, Spatial4D-Bench provides a more comprehensive evaluation benchmarks with 18 tasks. Since there are some tasks overlapping with existing benchmarks, it would be interesting to investigate the challenge of these tasks between existing benchmarks and Spatial4D-Bench. To this end, we compare with VSI-Bench [42], one of the most representative spatial intelligence benchmarks, on overlapping tasks using Qwen3-VL-30B-A3B, InternVL3.5-38B, and InternVL3.5-8B. As shown in Table˜3, overall, state-of-the-art models perform worse on our Spatial4D-Bench. Specifically, on some perception-related tasks, e.g., object size and room size, the performance is comparable across both benchmarks. However, for more challenging reasoning tasks, such as relative orientation and route plan, substantial divergences can be observed, especially on the route plan task, where the tested models exhibit a substantial performance drop of approximately 22% to 29% on Spatial4D-Bench compared to VSI-Bench [42]. In addition to the difficulty elevation described in Section˜3.3, this can be attributed to the fact that the route plan task in Spatial4D-Bench requires long-horizon planning and 4D spatial understanding over longer sequences and more complex state transitions, which distinguishes Spatial4D-Bench from existing benchmarks that may rely on shorter horizons or simpler topological graphs.
@@ 4.4 Further Analysis and Discussion
      Language Priors vs. Random Chance: In the absence of visual data, the model defaults to learned LLM priors, e.g., probabilistically associating “oven” with “kitchen”. As similarly observed on VSI-Bench [42], these priors often provide a statistical floor that is significantly better than random chance.
@@ 4.5 Qualitative Case Analysis
A major challenge in 4D spatial intelligence is spatiotemporal reasoning that requires understanding a coherent 4D world over time rather than treating videos as a collection of disjoint semantic concepts. By designing tasks that connect past, present, and future visual observations, Spatial4D-Bench reveals a significant temporal incoherence in SOTA MLLMs. Fig.˜6 shows failure and success cases from the action prediction task, which exposes this fragility. In Case 1 (failure), the ground truth action is “washing” a yogurt cup, but the model incorrectly predicts “placing it in the sink”, driven by a semantic prior rather than the specific temporal action. The benchmark’s design allows us to probe the cause via the model’s self-explanation: “…the person walks to the sink holding the empty cup… indicating they are about to place it in the sink rather than wash it or throw it away”. This textual feedback from the model demonstrates that the model successfully tracked the trajectory but failed to infer the latent intent (washing for recycling), proving that even powerful models fail to ground their predictions in the actual temporal dynamics when those dynamics conflict with training priors. On the other hand, Case 2 (success) validates the benchmark’s ability to measure robust chain-of-thought reasoning when explicit cues (OCR) are present. The model correctly infers the drink will taste “sweet” by synthesizing the label “sugar” with the stirring action (“mixing sugar into the drink makes it sweet…”). This contrast underscores the diagnostic value of our Spatial4D-Bench: it can distinguish between scenarios where models rely on robust textual grounding versus those where they collapse into hallucination due to temporal ambiguity.
@@
By incorporating the physical plausibility reasoning in the spatiotemporal reasoning category, specifically using AI-generated physical anomalies, Spatial4D-Bench provides a unique evaluation of the “intuitive physics engine” of MLLMs. Our evaluation exposes a sharp dissociation between low-level perceptual grounding and high-level physical knowledge. In Fig.˜7 Case 1 (failure), the benchmark challenges the model with a subtle violation of fluid dynamics. GPT-5 fails to identify a physical anomaly where liquid levels behave inconsistently during pouring, yet its self-explanation recites perfect high-level physics principles: “Matter doesn’t disappear from one place and reappear somewhere else… Teleportation breaks locality and conservation of mass…”. This result highlights a critical insight enabled by our benchmark: SOTA models possess abstract knowledge of physical laws but lack the visual grounding to detect their violation in pixel space. The model “knows” the laws of physics but cannot “see” them being broken, relying instead on analyzing the textual plausibility of the options (e.g., rejecting “teleportation” as a concept) rather than verifying the visual dynamics. Spatial4D-Bench thus serves as a necessary filter to differentiate between models that merely know physics textually and those that can perceive physics visually. In contrast, Case 2 (success) shows that the model is capable of detecting violations in rigid body dynamics. It correctly identifies that a baseball and bat exhibit unnatural softness and defy gravity, suggesting that conspicuous deviations in material properties and trajectory dynamics are easier for current architectures to flag than subtle fluid inconsistencies.
@@
Through evaluation on Spatial4D-Bench, we observe that perceptual ambiguity can lead to confident hallucinations, even in tasks (such as spatial memory) that require explicit tracking. In Fig.˜8 Case 1 (failure), the model incorrectly localizes the final position of a teapot. Crucially, the self-explanation reveals a confident hallucination: “After he pours the water… he places the teapot on the back-right burner (the burner directly above the original position)”. This contradicts the visual evidence of the teapot’s placement in other locations. This suggests that texture similarity across the stove top surface causes the model to lose track of the object’s specific geometric coordinates. Unlike a tracking failure where a model might express uncertainty, here the model constructs a coherent (but false) narrative to fill the perceptual gap. This failure mode validates the necessity of Spatial4D-Bench’s fine-grained annotation: unlike simpler existence or classification tasks, our spatial memory queries force the model to confront texture confusion and occlusion. The benchmark demonstrates that even when a model is confident and generates plausible-sounding narratives, it often lacks the precise metric grounding required for 4D spatial intelligence. By contrast, Case 2 (success) demonstrates that the model maintains robust temporal tracking when visual events are semantically distinct. The model correctly counts that the faucet was used “two” times. Its self-explanation, “…first to wash their hands at the sink, and later to fill a glass with water”, shows a successful linking of two separate temporal events, implying that the “grounding gap” is highly sensitive to visual saliency. The model fails when tracking requires resolving low-level texture ambiguity (Case 1) but succeeds when tracking high-level, semantically distinct actions (Case 2), further validating the benchmark’s ability to probe the granular limits of spatial memory.
@@
The route plan task in the spatial reasoning category illustrates the inability of current MLLMs to construct accurate mental maps from egocentric video streams. In Fig.˜9, the model attempts to plan a path from a hallway to a bathroom. While GPT-5 outputs a confident self-explanation, reasoning that a right turn at step 7 would “line up with the bedroom doorway”, this contradicts the visual geometry: the robot has just entered a hallway configuration where a left turn (option A) is geometrically required to face the bedroom entrance. The model hallucinates a spatial layout that fits its internal narrative but ignores the visual reality of the scene. Furthermore, while the model correctly identifies the final turn (step 11) as a right turn, this success is coincidental, derived from a flawed intermediate trajectory. This failure highlights a critical bottleneck: SOTA MLLMs struggle with long-horizon spatial consistency and relative orientation. By enforcing strict directional accuracy over multi-step paths, Spatial4D-Bench effectively disentangles true embodied route planning capabilities from lucky guesses driven by language priors.